# TREAD: Token Routing for Efficient Architecture-agnostic Diffusion Training

### v2

@@ Acknowledgement
This project has been supported by the German Federal Ministry for Economic Affairs and Climate Action within the project “NXT GEN AI METHODS – Generative Methoden für Perzeption, Prädiktion und Planung”, the bidt project KLIMA-MEMES, Bayer AG, and the German Research Foundation (DFG) project 421703927. The authors gratefully acknowledge the Gauss Center for Supercomputing for providing compute through the NIC on JUWELS at JSC and the HPC resources supplied by the Erlangen National High Performance Computing Center (NHR@FAU funded by DFG project 440719683) under the NHR project JA-22883.
Supplementary Material
@@ A.3 Hyperparameters
DiT-S | DiT-B | DiT-L | DiT-XL | |
Optimization | ||||
Batch size | 256 | 256 | 256 | 256 |
Optimizer | AdamW | AdamW | AdamW | AdamW |
LR | 1e-4 | 1e-4 | 1e-4 | 1e-4 |
(0.9, 0.999) | (0.9, 0.999) | (0.9, 0.999) | (0.9, 0.999) | |
Optimization - Finetune | ||||
Batch size | - | - | - | 1,024 |
Optimizer | - | - | - | AdamW |
LR | - | - | - | 1e-5 |
- | - | - | (0.9, 0.999) | |
Architecture | ||||
Dim | 384 | 768 | 1,024 | 1,152 |
Heads | 6 | 12 | 16 | 16 |
Layers | 12 | 12 | 24 | 28 |
TREAD | ||||
Route | ||||
Selection Rate | 0.5 | 0.5 | 0.5 | 0.5 |

### v3

@@ Acknowledgement
This project has been supported by the Federal Ministry for Economic Affairs and Energy (BMWE) within the project “NXT GEN AI METHODS - Generative Methoden für Perzeption, Prädiktion und Planung”, the project “GeniusRobot” (01IS24083) funded by the Federal Ministry of Research, Technology and Space (BMFTR), and the bidt project KLIMA-MEMES, and partially funded by the Horizon Europe project ELLIOT (101214398). The authors gratefully acknowledge the Gauss Center for Supercomputing for providing compute through the NIC on JUWELS/JUPITER at JSC and the HPC resources supplied by the NHR @ FAU Erlangen.
@@ A.3 Hyperparameters
DiT-S DiT-B DiT-L DiT-XL Optimization Batch size 256 256 256 256 Optimizer AdamW AdamW AdamW AdamW LR 1e-4 1e-4 1e-4 1e-4 (0.9, 0.999) (0.9, 0.999) (0.9, 0.999) (0.9, 0.999) Optimization - Finetune Batch size - - - 1,024 Optimizer - - - AdamW LR - - - 1e-5 - - - (0.9, 0.999) Architecture Dim 384 768 1,024 1,152 Heads 6 12 16 16 Layers 12 12 24 28 TREAD Route Selection Rate 0.5 0.5 0.5 0.5
@@ A.4 Classifier-Free Guidance
Appendix B Loss Curves and Routing induced Loss Gap
We provide loss curves in Figure S10 for better understanding of the interaction between loss, route length and FID. The loss gap to the baseline DiT can be explained using the increased difficulty during training which is induced by longer routes.